A SAMPLE PAPER

WEBMASTER

One way to make your paper available to others is to post a dvi version of the file—just like the one are looking at. This works well if your paper does not contain figures or non-standard fonts. Other options include:

1. posting TEX source files (this requires the receipient to run TEX),

2. posting the Postscript file (the drawback of this method is that Postscript files tend to be rather large),

3. converting the file to the pdf format (this is not yet available on our system, but the installation is planned), and

4. converting the file to the html format using latex2html (the version currently available on our system does not work well—an upgrade to a newer version is planned, but, judging from the current version, this method is suitable only for small and simple documents)

When posting files through your home page, remember to put them into your public html directory, make them readable by everyone, and give full and correct URLs (e.g. http://math.lsu.edu/∼your login name/paper.dvi).